OpenStudy (anonymous):

Solve the exponential equation. Express your solutions in exact form only. Please show all of your work. 4^x-10*2^x-74=1

OpenStudy (anonymous):

\[(2^{x})^{2}-10\times2^{x}-78=0\] let \[2^x=t\],t>0 then \[t^2-10t-78=0\] then solve this equation and leave out the negative value t=\[5+\sqrt{103}\] so x=\[\log_{2} (5+\sqrt{103})\]

OpenStudy (anonymous):

4^x-10*2^x-74=1 4^x-10*2^x-75=0 (2)^2x -10(2)^x -75=0 (2^x)^2 -10(2^x) -75=0 m=2^x so that m^2 -10m -75=0, factoring or solved by quadratic formula (m+5)(m-15)=0, m=-5 and m = 15 now using m=2^x for m=15, 2^x=15 taking log on both sides log2^x=log15 xlog2=log15 x=log15/ log2 ans....... if using m=-5=2^x taking log on both sides log(-5)=log2^x therefore we cant use -5

OpenStudy (anonymous):

which one of you is correct?